# Synecdoche: Efficient and Accurate In-Network Traffic Classification via Direct Packet Sequential Pattern Matching

### v1

@@ Abstract
Traffic classification on programmable data plane holds great promise for line-rate processing, with methods evolving from per-packet to flow-level analysis for higher accuracy. However, a trade-off between accuracy and efficiency persists. Statistical feature-based methods align with hardware constraints but often exhibit limited accuracy, while online deep learning methods using packet sequential features achieve superior accuracy but require substantial computational resources. This paper presents Synecdoche, the first traffic classification framework that successfully deploys packet sequential features on a programmable data plane via pattern matching, achieving both high accuracy and efficiency. Our key insight is that discriminative information concentrates in short sub-sequences—termed Key Segments—that serve as compact traffic features for efficient data plane matching. Synecdoche employs an ”offline discovery, online matching” paradigm: deep learning models automatically discover Key Segment patterns offline, which are then compiled into optimized table entries for direct data plane matching. Extensive experiments demonstrate Synecdoche’s superior accuracy, improving F1-scores by up to 26.4% against statistical methods and 18.3% against online deep learning methods, while reducing latency by 13.0% and achieving 79.2% reduction in SRAM usage. The source code of Synecdoche [^1] https://github.com/swampx/Synecdoche is publicly available to facilitate reproducibility and further research.
@@ VI-A Experimental Setup
  • Data Plane Methods:
    NetBeacon [7]: State-of-the-art statistical feature-based approach using decision trees on programmable switches, representing traditional ML-based P4 inference. Brain-on-Switch [15]: A representative framework for deploying quantized neural networks onto the data plane.
  • Offline Baselines(Upper Bound Analysis):
    Random Forest: Offline statistical method using the same features as NetBeacon but without hardware constraints (unlimited tree depth and estimators). FS-Net [31]: State-of-the-art offline sequence feature-based deep learning method, representing the upper bound of accuracy for sequence-based approaches.
To ensure fair comparison, we configure the input features and hyperparameters of comparing methods according to the settings specified in their respective official implementations[37, 38, 36]. Notably, regarding the IMIS system in BoS, while we recognize its contribution in improving classification performance through data-control plane interaction, our evaluation focuses on the classification capability of the data plane itself, without incorporating the performance gains from such interaction. This methodology ensures fairness, as Synecdoche can also be integrated into systems like IMIS.
@@ VI-D Performance and Efficiency Analysis (RQ2)
Table IV presents comprehensive performance metrics. In terms of per-flow storage requirements, our approach maintains the feature information of the most recent $L_{max}$ packets, including: a 32-bit direction register storing the source IP of the first packet for subsequent direction determination, a 32-bit register for the latest packet timestamp, an 8-bit counter for the current packet count, and $L_{max}\times 16$ bits for packet features. When $L_{max}=4$, each flow requires 136 bits of storage., significantly outperforming both baselines. NetBeacon stores different statistical features totaling 272 bits per flow, while BoS requires 288 bits to maintain sliding window hidden states (window size=8) plus per-class cumulative scores for classification decisions. Both Synecdoche variants demonstrate superior resource efficiency in terms of SRAM utilization. The SRAM version achieves exceptional memory efficiency with only 4.03% SRAM usage, representing a remarkable 79.2% reduction compared to NetBeacon (19.38%) and an 85.8% reduction compared to BoS (28.33%).
@@ VII Conclusion
This paper presents Synecdoche, a novel traffic classification framework that successfully bridges the accuracy-efficiency gap on programmable data plane through direct packet sequential pattern matching. By leveraging an ”offline discovery, online matching” paradigm that automatically extracts Key Segments using deep learning models and deploys them as optimized table entries, Synecdoche significantly outperforms existing statistical and deep learning approaches while substantially reducing hardware resource consumption and processing latency. This work represents a fundamental advance in programmable data plane traffic classification, demonstrating that intelligent pattern extraction and hardware-optimized deployment strategies can overcome traditional accuracy-efficiency trade-offs and pave the way for sophisticated packet sequence analysis at line rate in emerging network applications. Our future work will continue to explore the impact of network fluctuations on Key Segment effectiveness and investigate mechanisms for efficient rule updates in deployed systems.
@@ Acknowledgment
This work was supported by The Cybersecurity Professional Committee of the Chinese Association of Educational Technology(NO. 2023CAET1002)

### v2

@@ Abstract
Traffic classification on programmable data plane holds great promise for line-rate processing, with methods evolving from per-packet to flow-level analysis for higher accuracy. However, a trade-off between accuracy and efficiency persists. Statistical feature-based methods align with hardware constraints but often exhibit limited accuracy, while online deep learning methods using packet sequential features achieve superior accuracy but require substantial computational resources. This paper presents Synecdoche, the first traffic classification framework that successfully deploys packet sequential features on a programmable data plane via pattern matching, achieving both high accuracy and efficiency. Our key insight is that discriminative information concentrates in short sub-sequences—termed Key Segments—that serve as compact traffic features for efficient data plane matching. Synecdoche employs an ”offline discovery, online matching” paradigm: deep learning models automatically discover Key Segment patterns offline, which are then compiled into optimized table entries for direct data plane matching. Extensive experiments demonstrate Synecdoche’s superior accuracy, improving F1-scores by up to 26.4% against statistical methods and 18.3% against online deep learning methods, while reducing latency by 13.0% and achieving 79.2% reduction in SRAM usage.
@@ VI-A Experimental Setup
  • Data Plane Methods: NetBeacon [7]: State-of-the-art statistical feature-based approach using decision trees on programmable switches, representing traditional ML-based P4 inference. Brain-on-Switch [15]: A representative framework for deploying quantized neural networks onto the data plane.
  • Offline Baselines (Upper Bound Analysis): Random Forest: Offline statistical method using the same features as NetBeacon but without hardware constraints (unlimited tree depth and estimators). FS-Net [31]: State-of-the-art offline sequence feature-based deep learning method, representing the upper bound of accuracy for sequence-based approaches.
To ensure fair comparison, we configure the input features and hyperparameters of comparing methods according to the settings specified in their respective official implementations[37, 38, 36]. For the IMIS system in BoS, we evaluate only the data plane’s classification capability, excluding gains from data-control plane interaction, since Synecdoche can likewise be integrated into such systems.
@@ VI-D Performance and Efficiency Analysis (RQ2)
Table IV presents comprehensive performance metrics. In terms of per-flow storage requirements, our approach maintains the feature information of the most recent $L_{max}$ packets, including: a 32-bit direction register storing the source IP of the first packet for subsequent direction determination, a 32-bit register for the latest packet timestamp, an 8-bit counter for the current packet count, and $L_{max}\times 16$ bits for packet features. When $L_{max}=4$, each flow requires 136 bits of storage, significantly outperforming both baselines. NetBeacon stores different statistical features totaling 272 bits per flow, while BoS requires 288 bits to maintain sliding window hidden states (window size=8) plus per-class cumulative scores for classification decisions. Both Synecdoche variants demonstrate superior resource efficiency in terms of SRAM utilization. The SRAM version achieves exceptional memory efficiency with only 4.03% SRAM usage, representing a remarkable 79.2% reduction compared to NetBeacon (19.38%) and an 85.8% reduction compared to BoS (28.33%).
@@ VII Conclusion
This paper presents Synecdoche, a novel traffic classification framework that successfully bridges the accuracy-efficiency gap on programmable data plane through direct packet sequential pattern matching. By leveraging an ”offline discovery, online matching” paradigm that automatically extracts Key Segments using deep learning models and deploys them as optimized table entries, Synecdoche significantly outperforms existing statistical and deep learning approaches while substantially reducing hardware resource consumption and processing latency. This work represents a fundamental advance in programmable data plane traffic classification, demonstrating that intelligent pattern extraction and hardware-optimized deployment strategies can overcome traditional accuracy-efficiency trade-offs and pave the way for sophisticated packet sequence analysis at line rate in emerging network applications. Our future work will continue to explore the impact of network fluctuations on Key Segment effectiveness and investigate mechanisms for efficient rule updates in deployed systems. The authors have provided public access to their code at https://github.com/swampx/Synecdoche.
@@ Acknowledgment
This work was partly supported by The Cybersecurity Professional Committee of the Chinese Association of Educational Technology(NO. 2023CAET1002)